OpenStudy (anonymous):

I WILL BECOME A FAN AND GIVE A MEDAL What is the value of c so that y = x^2 + 10x + c is a perfect square trinomial? A. 10 B. 25 C. 50 D. 100 Please tell me the answer and how you got it! Thanks ya'all!

OpenStudy (anonymous):

B, u have to divide 10 by to then square it

OpenStudy (anonymous):

becasuse 10 is the 2nd coefficient

OpenStudy (anonymous):

Thanks so much! Could you help me with another one please? :)

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

What is the value of c so that x^2+9x+c is a perfect square trinomial? A. 81 B. 81/2 C. 81/4 D. 9/2 Thanks again! :)

OpenStudy (zzr0ck3r):

is half the b term squared

OpenStudy (anonymous):

C, 4.5^2

OpenStudy (the_fizicx99):

\(\ (a+b)^2 = a^2 + 2(a)(b) + b^2 \) A is x and B would be 5 because \(\ \dfrac{10}{2} = 5 \) \(\ x^2 + 2(x)(5) + 5^2 = x^2 + 10x + 25 \)